# Collapse-based models for gravity do not violate the entanglement-based witness of non-classicality

### v2

@@ I Introduction
A particularly promising approach to testing quantum gravity has recently been proposed, based on a novel “witness of non-classicality”. This witness relies on the entangling power of a given system to conclude that the system has non-classical features. In particular, these tests are based on the so-called General Witness Theorem (GWT), [cmvvRevModPhys.97.015006, marletto_witnessing_2020], stating that if a system $M$ (such as gravity) can mediate (by local means) entanglement between two quantum systems, $A$ and $B$, (e.g. two masses) then it must be non-classical, [marletto_witnessing_2020]. By “local means” here we mean a specific protocol, detailed in [marletto_witnessing_2020, bose_spin_2017, marletto_gravitationally_2017], where $A$ and $B$ must not interact directly with each other, but only via the mediator $M$, as schematically represented in Fig.1 (a). Interestingly, “non-classicality” is a theory-independent generalisation of what in quantum theory is expressed as “having at least two distinct physical variables that do not commute”, which can be expressed within a general information-theoretic framework, the constructor theory of information, [deutsch_constructor_2015]. Informally, being non-classical means having two or more distinct physical variables that cannot simultaneously be measured to an arbitrarily high degree of accuracy [marletto_witnessing_2020].
Due to the generality of GWT, it offers a broad theoretical basis for recently proposed experiments that can test quantum effects in gravity at the laboratory scale, based on the generation of gravitational entanglement between two massive probes (see Fig. 1 (b)) - the so-called Bose-Marletto-Vedral effect, [bose_spin_2017, marletto_gravitationally_2017]. It also provides a basis for any other experiment that (beyond the case of gravity) intends to show that some system $M$ is non-classical [raia_role_2024], using the effect of its entangling power.
A particularly appealing feature of the GWT is that, by using the constructor theory of information, it avoids assuming the usual machinery of quantum information theory, thus extending beyond quantum theory existing results, such as the theorems that forbid the creation of entanglement via local operations and classical communication. Moreover, the GWT is proven without assuming the existence of a probability space, in contrast to existing approaches such as Generalised Probabilistic Theories [plavala_general_2023]. This generality is particularly important as one wants to use it in a context where the system $M$ may or may not obey quantum theory itself.
@@
Recently, some claims have been made that the collapse-based model of classical gravity, i.e. Diósi-Penrose model [penrose1996gravity, diosi1987universal, DiosiPhysRevA.40.1165, Tilloy_2016], can predict gravitationally induced entanglement between quantum objects [trillo2024diosipenrosemodelclassicalgravity, angeli2025positivityentanglementmarkovianopen], resulting in gravitationally induced entanglement is insufficient to conclude that gravity is fundamentally quantum [trillo2024diosipenrosemodelclassicalgravity]. These claims are contrary to the GWT statement. Here we vindicate the GWT and show that collapse-based models of gravity are fundamentally nonlocal, violating the assumption of locality. The generation of entanglement between masses stems from the hidden quantum-like monitoring process, not from the backaction of the gravitational field. That is, the gravitational field does not induce the entanglement. Even though the interaction between quantized matter and the gravitational field cannot generate entanglement, the backaction of the gravitational field is nonlocal itself. Besides, the monitoring process as a part of the dynamics violates the principle of locality and reveals quantumlike features.
@@ II Brief Review of collapse-based models of gravity
Here we briefly review the collapse-based models of gravity [diosi1987universal, DiosiPhysRevA.40.1165, penrose1996gravity, Bahrami_2014]. Through the lens of continuous measurement theory, Tilloy and Diósi present a linear version of collapse-based models, i.e. DP model [Tilloy_2016].
@@
where $\hat{{\varrho}}_{\sigma}(\mathbf{r})=\sum_{n=1}^{N}m_{n}g_{\sigma}(\mathbf{r}-\hat{\mathbf{x}}_{n})$ (NN is the particle number) representing density of matter and $\delta{\varrho}_{t}(\mathbf{r})$ denotes the white noise in time, satisfying $\mathbb{E}[\delta{\varrho}_{t}(\mathbf{r})\delta{\varrho}_{\tau}(\mathbf{s})]=\gamma^{-1}_{\mathbf{r}\mathbf{s}}\delta(t-\tau)$. Note that the function $g_{\sigma}$ is a normalized Gaussian of width $\sigma$ [Tilloy_2016]. Here, $\gamma_{rs}$ is a non-negative kernel that intuitively encodes the correlation between the detectors at positions $r$ and $s$. After monitoring by a hidden detector, the stochastic master equation (SME) that governs the dynamics of the density matrix of the system is given by [Tilloy_2016]:
@@
where $\mathcal{H}[{\hat{\varrho}}_{\sigma}(\mathbf{r})](\hat{\rho})=\left\{{\hat{\varrho}}_{\sigma}(\mathbf{r})-\left<{\hat{\varrho}}_{\sigma}(\mathbf{r})\right>_{t},{\hat{\rho}}_{t}\right\}$ and $(\sigma)$ denotes an optional convolution with $g_{\sigma}$. This is the standard SME if one replaces $\int dr{\hat{\rho}}_{\sigma}(r)$ as another operator $\hat{A}$ [wiseman2009quantum]. Note that we have set $\hbar=1$. If we average the measurement outcome of the hidden detector, the state satisfies
@@
The modified semiclassical Newton potential now becomes stochastic and takes the form $\Phi(\mathbf{r})=-G\int\mathrm{d}\mathbf{s}\frac{{\varrho}(\mathbf{s})}{|\mathbf{r}-\mathbf{s}|}$ [Tilloy_2016]. That is, the semiclassical interaction ${\hat{V}}_{G}$ is now given as
@@
where $\mathrm{d}{\hat{\rho}}^{\text{SEM}}$ is the change of quantum state according to Eq. (II) in the monitoring procese. Expanding the exponential up to second order in Eq. (6) then gives the average SEM for the complete evolution[Tilloy_2016]:
@@
where ${\hat{V}}_{G,\sigma}=-\frac{G}{2}\int\mathrm{d}\mathbf{r}\mathrm{d}\mathbf{s}\frac{{\hat{\varrho}}_{\sigma}(\mathbf{r}){\hat{\varrho}}_{(\sigma)}(\mathbf{s})}{|\mathbf{r}-\mathbf{s}|}$ is an effective potential term. The CSL and DP models both adhere to the stochastic master equation of Eq. (II), with the primary distinction between them being the choice of the nontrivial correlator $\gamma_{\mathbf{r}\mathbf{s}}$[Tilloy_2016]. For the CSL model, the correlator is chosen as $\gamma_{\mathbf{r}\mathbf{s}}=\gamma\delta(\mathbf{r}-\mathbf{s})$. For DP model, $\gamma_{\mathbf{r}\mathbf{s}}=\kappa G\frac{1}{|\mathbf{r}-\mathbf{s}|}$ or $\gamma_{\mathbf{r}\mathbf{s}}^{-1}=-\frac{1}{4\pi\kappa G}\delta(t-\tau)\nabla^{2}\delta(\mathbf{r}-\mathbf{s})$.
@@ III Entanglement generation in the DP model
For analysing the dynamics of the DP model in detail, we set $\gamma_{\mathbf{r}\mathbf{s}}=2G\frac{1}{|\mathbf{r}-\mathbf{s}|}$ and substitute it into Eq. II, resulting evolution of the equation of the DP model,
@@
This equation has the same form as in [trillo2024diosipenrosemodelclassicalgravity].
@@
with the corresponding density matrix: $\rho(0)=\frac{1}{4}\sum_{x_{1},x_{2}}|x_{1}\rangle\langle x_{1}|\otimes|x_{2}\rangle\langle x_{2}|,$ where $x_{i}\in\{0,1\}$. In [trillo2024diosipenrosemodelclassicalgravity, angeli2025positivityentanglementmarkovianopen], the authors show that the collapsed-based models of gravity, i.e. DP model, can predict entanglement between masses 1 and 2, which may conflict with GWT.
@@ IV What goes wrong in the argument
Recently, Trillo and Navascué claim the gravitationally induced entanglement is insufficient to conclude that gravity is fundamentally quantum since the DP model can generate gravitationally induced entanglement between two quantum masses [trillo2024diosipenrosemodelclassicalgravity]. Here we disprove their claims and show that their work does not violate the entanglement-based witness of non-classicality, e.g., GWT. In the DP model [Tilloy_2016, trillo2024diosipenrosemodelclassicalgravity], the generation of entanglement between two masses is not induced by classical gravity but instead by a hidden monitoring process with quantum features. The DP model is fundamentally nonlocal or has a quantum degree of freedom to mediate the interaction between matter. Thus, the claim that “the Diósi-Penrose model of classical gravity predicts gravitationally induced entanglement” is inaccurate and not true.
@@
On the other hand, from the quantum information perspective, if the gravitational field is classical (scalar function $\Phi(\mathbf{r})$ and it must have only one basis $\{|{i}\rangle\langle{i}|\}$). As an entity, the purely classical gravitational field can not mediate entanglement between two quantum masses [marletto_gravitationally_2017, cmvvRevModPhys.97.015006]. One more point needs to be clarified: even though the back-action of the gravitational field in the DP model can not generate entanglement, it is fundamentally nonlocal.
2. About the entanglement generation. The entanglement between masses can be generated in the DP model [trillo2024diosipenrosemodelclassicalgravity, angeli2025positivityentanglementmarkovianopen], but it is not induced by gravity. The nonrival hidden detector induces the entanglement of the DP model [Tilloy_2016, trillo2024diosipenrosemodelclassicalgravity, angeli2025positivityentanglementmarkovianopen] with the non-trivial choice of $\gamma_{\mathbf{r}\mathbf{s}}=2G\frac{1}{|\mathbf{r}-\mathbf{s}|}$. This correlator is nonlocal, as mentioned in other nonlinear gravity works [Bahrami_2014].
@@
where $m_{1}=m_{2}$ and $g_{\sigma}(\mathbf{r}-\mathbf{x}_{i})=\frac{e^{-|\mathbf{r}-\mathbf{x}_{i}|^{2}/(2\sigma)}}{(2\pi\sigma)^{3/2}}$. Since $I_{\mathbf{x}_{i},\mathbf{x}_{j}}=\int drds\frac{1}{|{\mathbf{r}-\mathbf{s}}|}g_{\sigma}(\mathbf{r}-\mathbf{x}_{i})g_{\sigma}(\mathbf{s}-\mathbf{x}_{j})=\frac{\text{erf}(|\mathbf{x}_{i}-\mathbf{x}_{j}|/(2\sigma))}{|\mathbf{x}_{i}-\mathbf{x}_{j}|}\quad\text{or}\quad\frac{1}{\sigma\sqrt{\pi}}$ for $|\mathbf{x}_{i}-\mathbf{x}_{j}|\neq 0$ and $|\mathbf{x}_{i}-\mathbf{x}_{j}|=0$ respectively [trillo2024diosipenrosemodelclassicalgravity]. Remarkably, $\text{erf}(x)$ is the error function. Without loss of generality, one may set $\tilde{f}(z)=\text{erf}(z/2\sigma)/z$ if $z\neq 0$ and $\tilde{f}(z)=\frac{1}{\sigma\sqrt{\pi}}$ if $z=0$. As shown in Figure 1 (b), in the BMV setup, two particles are on the same horizontal line, so we can replace $\mathbf{x}$ with the scalar $x$. One can directly compute the density matrix element as:
@@
where $x_{1},x_{2},y_{1},y_{2}\in\{0,1\}$ representing the position of the corresponding particle. Since $\mathbf{x}_{1}=\pm\frac{a}{2}$ ($|{0}\rangle_{1}:=|{x_{1}=-\frac{a}{2}}\rangle$ and $|{1}\rangle_{1}:=|{x_{1}=+\frac{a}{2}}\rangle$), and $\mathbf{x}_{2}=\pm\frac{a}{2}+d$ ($|{0}\rangle_{2}:=|{x_{2}=d-\frac{a}{2}}\rangle$ and $|{1}\rangle_{2}:=|{x_{2}=d+\frac{a}{2}}\rangle$), the negativity of the density matrix ${\hat{\rho}}(\Delta t)$ is calculated as
$\mathcal{N}({\hat{\rho}}^{T1}(\Delta t))=\text{Max}(0,p)+\text{Max}(0,q),$ (14)
where $p=\frac{Gm^{2}\Delta t}{8\hbar}(2\tilde{f}(a)+\tilde{f}(d)+\tilde{f}(d+2a)-2\tilde{f}(0)-2\tilde{f}(a+d))$ and $q=\frac{Gm^{2}\Delta t}{8\hbar}(2\tilde{f}(a)+2\tilde{f}(d+a)-2\tilde{f}(0)-\tilde{f}(d)-\tilde{f}(2a+d))$. Note that here we write $\hbar$ back. Since $\tilde{f}(z)=1/(\sigma\sqrt{\pi})+\frac{z^{2}}{12\sqrt{\pi}\sigma^{3}}+O(z^{3}/\sigma^{4})$, $\tilde{f}$ is a convex function and $\tilde{f}(z)>0$ ($z>0$), thus $q>0$. One can verify that $q=0+O(z^{3}/\sigma^{4})$ resulting
[…]
The above analysis indicates $\sigma\gg\{a,d,d+a,2a+d\}$ so that one can omit the high order term $O(z^{3}/\sigma^{4})$ in $\tilde{f}(z)$. This suggests that entanglement indeed can be generated in the monitoring process by Eq. (II).
3. About the quantum hidden degree of freedom in the monitoring process. In the collapse-based model of gravity [trillo2024diosipenrosemodelclassicalgravity, Tilloy_2016, angeli2025positivityentanglementmarkovianopen], the SEM (see Eq. II) is used to analyze the DP and CML models. This SME stems from the continuous measurement theory in which there is a hidden degree of freedom that can monitor the observable of interest [Tilloy_2016]. In the DP case, this hidden degree of freedom (also called ‘hidden detector’) weakly measures the mass density. It is well known that no consistent quantum-classical dynamics for quantum measurement problems [diosi2000quantum, diosi2014hybrid]. Solving the measurement problem either requires introducing non-commutative degrees of freedom into classical degrees of freedom (which contradicts the definition of classical systems) or requires the introduction of superselection rules [sudarshan1976interaction, koopman1931hamiltonian]. Indeed, continuous measurement theory is also introduced to address the quantum measurement problem, and it introduces a quantum hidden detector [wiseman2009quantum]. Thus, this hidden degree of freedom must be quantum-like. If the hidden degree of freedom is classical, it can not be used to monitor the observables of quantum systems. In [Tilloy_2016], the author made the following statement:
@@
If the monitoring process of hidden detectors is not a physical process but merely a mathematical construct [Tilloy_2016, trillo2024diosipenrosemodelclassicalgravity, angeli2025positivityentanglementmarkovianopen], collapse-based models would violate causality and the principle of locality, as they implicitly assume instantaneous interactions between masses. This assumption further leads to the paradoxical conclusion that all matter would spontaneously decohere and (weakly) entangle together.

### v3

@@ I Introduction
A particularly promising approach to testing quantum gravity has recently been proposed, based on a novel “witness of non-classicality”. This witness relies on the entangling power of a given system to conclude that the system has non-classical features. In particular, these tests are based on the so-called General Witness Theorem (GWT), [19, 17], stating that if a system $M$ (such as gravity) can mediate (by local means) entanglement between two quantum systems, $A$ and $B$, (e.g. two masses) then it must be non-classical, [17]. By “local means” here we mean a specific protocol, detailed in [17, 4, 15], where $A$ and $B$ must not interact directly with each other, but only via the mediator $M$, as schematically represented in Fig.1 (a). Interestingly, “non-classicality” is a theory-independent generalisation of what in quantum theory is expressed as “having at least two distinct physical variables that do not commute”, which can be expressed within a general information-theoretic framework, the constructor theory of information, [5]. Informally, being non-classical means having two or more distinct physical variables that cannot simultaneously be measured to an arbitrarily high degree of accuracy [17].
Due to the generality of GWT, it offers a broad theoretical basis for recently proposed experiments that can test quantum effects in gravity at the laboratory scale, based on the generation of gravitational entanglement between two massive probes (see Fig. 1 (b)) - the so-called Bose-Marletto-Vedral effect, [4, 15]. It also provides a basis for any other experiment that (beyond the case of gravity) intends to show that some system $M$ is non-classical [23], using the effect of its entangling power.
A particularly appealing feature of the GWT is that, by using the constructor theory of information, it avoids assuming the usual machinery of quantum information theory, thus extending beyond quantum theory existing results, such as the theorems that forbid the creation of entanglement via local operations and classical communication. Moreover, the GWT is proven without assuming the existence of a probability space, in contrast to existing approaches such as Generalised Probabilistic Theories [22]. This generality is particularly important as one wants to use it in a context where the system $M$ may or may not obey quantum theory itself.
@@
Recently, some claims have been made that the collapse-based model of classical gravity [21, 8, 6, 25], i.e. Diósi-Penrose model, can predict gravitationally induced entanglement between quantum objects [26, 1, 13, 14], resulting in gravitationally induced entanglement is insufficient to conclude that gravity is fundamentally quantum [26]. These claims are contrary to the GWT statement. Here we vindicate the GWT and show that collapse-based models of gravity are fundamentally nonlocal, violating the assumption of locality. The generation of entanglement between masses stems from the hidden quantum-like monitoring process, not from the backaction of the gravitational field. That is, the gravitational field does not induce the entanglement. Even though the interaction between quantized matter and the gravitational field cannot generate entanglement, the backaction of the gravitational field is nonlocal itself. Besides, the monitoring process as a part of the dynamics violates the principle of locality and reveals quantum-like features.
@@ II Brief Review of collapse-based models of gravity
Here we briefly review the collapse-based models of gravity [8, 6, 21, 3]. Through the lens of continuous measurement theory, Tilloy and Diósi present a linear version of collapse-based models, i.e. DP model [25].
@@
where $\hat{{\varrho}}_{\sigma}(\mathbf{r})=\sum_{n=1}^{N}m_{n}g_{\sigma}(\mathbf{r}-\hat{\mathbf{x}}_{n})$ (NN is the particle number) representing density of matter and $\delta{\varrho}_{t}(\mathbf{r})$ denotes the white noise in time, satisfying $\mathbb{E}[\delta{\varrho}_{t}(\mathbf{r})\delta{\varrho}_{\tau}(\mathbf{s})]=\gamma^{-1}_{\mathbf{r}\mathbf{s}}\delta(t-\tau)$. Note that the function $g_{\sigma}$ is a normalized Gaussian of width $\sigma$ [25]. Here, $\gamma_{rs}$ is a non-negative kernel that intuitively encodes the correlation between the detectors at positions $r$ and $s$. After monitoring by a hidden detector, the stochastic master equation (SME) that governs the dynamics of the density matrix of the system is given by [25]:
@@
where $\mathcal{H}[{\hat{\varrho}}_{\sigma}(\mathbf{r})](\hat{\rho})=\left\{{\hat{\varrho}}_{\sigma}(\mathbf{r})-\left<{\hat{\varrho}}_{\sigma}(\mathbf{r})\right>_{t},{\hat{\rho}}_{t}\right\}$ and $(\sigma)$ denotes an optional convolution with $g_{\sigma}$. This is the standard SME if one replaces $\int dr{\hat{\rho}}_{\sigma}(r)$ as another operator $\hat{A}$ [27]. Note that we have set $\hbar=1$. If we average the measurement outcome of the hidden detector, the state satisfies
@@
The modified semiclassical Newton potential now becomes stochastic and takes the form $\Phi(\mathbf{r})=-G\int\mathrm{d}\mathbf{s}\frac{{\varrho}(\mathbf{s})}{|\mathbf{r}-\mathbf{s}|}$ [25]. That is, the semiclassical interaction ${\hat{V}}_{G}$ is now given as
@@
where $\mathrm{d}{\hat{\rho}}^{\text{SEM}}$ is the change of quantum state according to Eq. (II) in the monitoring procese. Expanding the exponential up to second order in Eq. (6) then gives the average SEM for the complete evolution[25]:
@@
where ${\hat{V}}_{G,\sigma}=-\frac{G}{2}\int\mathrm{d}\mathbf{r}\mathrm{d}\mathbf{s}\frac{{\hat{\varrho}}_{\sigma}(\mathbf{r}){\hat{\varrho}}_{(\sigma)}(\mathbf{s})}{|\mathbf{r}-\mathbf{s}|}$ is an effective potential term. The CSL and DP models both adhere to the stochastic master equation of Eq. (II), with the primary distinction between them being the choice of the nontrivial correlator $\gamma_{\mathbf{r}\mathbf{s}}$[25]. For the CSL model, the correlator is chosen as $\gamma_{\mathbf{r}\mathbf{s}}=\gamma\delta(\mathbf{r}-\mathbf{s})$. For DP model, $\gamma_{\mathbf{r}\mathbf{s}}=\kappa G\frac{1}{|\mathbf{r}-\mathbf{s}|}$ or $\gamma_{\mathbf{r}\mathbf{s}}^{-1}=-\frac{1}{4\pi\kappa G}\delta(t-\tau)\nabla^{2}\delta(\mathbf{r}-\mathbf{s})$.
@@ III Entanglement generation in the DP model
For analysing the dynamics of the DP model in detail, we set $\gamma_{\mathbf{r}\mathbf{s}}=2G\frac{1}{|\mathbf{r}-\mathbf{s}|}$ and substitute it into Eq. (II), resulting evolution of the equation of the DP model,
@@
This equation has the same form as in [26].
@@
with the corresponding density matrix: $\rho(0)=\frac{1}{4}\sum_{x_{1},x_{2}}|x_{1}\rangle\langle x_{1}|\otimes|x_{2}\rangle\langle x_{2}|,$ where $x_{i}\in\{0,1\}$. In [26, 1], the authors show that the collapsed-based models of gravity, i.e. DP model, can predict entanglement between masses 1 and 2, which may conflict with GWT.
@@ IV What goes wrong in the argument
Recently, Trillo and Navascué claim the gravitationally induced entanglement is insufficient to conclude that gravity is fundamentally quantum since the DP model can generate gravitationally induced entanglement between two quantum masses [26]. Here we disprove their claims and show that their work does not violate the entanglement-based witness of non-classicality, e.g., GWT. In the DP model [25, 26], the generation of entanglement between two masses is not induced by classical gravity but instead by a hidden monitoring process with quantum features. The DP model is fundamentally nonlocal or has a quantum degree of freedom to mediate the interaction between matter. Thus, the claim that “the Diósi-Penrose model of classical gravity predicts gravitationally induced entanglement” is inaccurate and not true.
@@
On the other hand, from the quantum information perspective, if the gravitational field is classical (scalar function $\Phi(\mathbf{r})$ and it must have only one basis $\{|{i}\rangle\langle{i}|\}$). As an entity, the purely classical gravitational field can not mediate entanglement between two quantum masses [15, 19]. One more point needs to be clarified: even though the back-action of the gravitational field in the DP model can not generate entanglement, it is fundamentally nonlocal.
2. About the entanglement generation. The entanglement between masses can be generated in the DP model [26, 1], but it is not induced by gravity. The nonrival hidden detector induces the entanglement of the DP model [25, 26, 1] with the non-trivial choice of $\gamma_{\mathbf{r}\mathbf{s}}=2G\frac{1}{|\mathbf{r}-\mathbf{s}|}$. This correlator is nonlocal, as mentioned in other nonlinear gravity works [3].
@@
where $m_{1}=m_{2}$ and $g_{\sigma}(\mathbf{r}-\mathbf{x}_{i})=\frac{e^{-|\mathbf{r}-\mathbf{x}_{i}|^{2}/(2\sigma)}}{(2\pi\sigma)^{3/2}}$. Note that $I_{\mathbf{x}_{i},\mathbf{x}_{j}}=\int drds\frac{1}{|{\mathbf{r}-\mathbf{s}}|}g_{\sigma}(\mathbf{r}-\mathbf{x}_{i})g_{\sigma}(\mathbf{s}-\mathbf{x}_{j})=\frac{\text{erf}(|\mathbf{x}_{i}-\mathbf{x}_{j}|/(2\sigma))}{|\mathbf{x}_{i}-\mathbf{x}_{j}|}\quad\text{or}\quad\frac{1}{\sigma\sqrt{\pi}}$ for $|\mathbf{x}_{i}-\mathbf{x}_{j}|\neq 0$ and $|\mathbf{x}_{i}-\mathbf{x}_{j}|=0$ respectively [26], where $\text{erf}(x)$ is the error function. Without loss of generality, one may set $\tilde{f}(z)=\text{erf}(z/2\sigma)/z$ if $z\neq 0$ and $\tilde{f}(z)=\frac{1}{\sigma\sqrt{\pi}}$ if $z=0$. As shown in Figure 1 (b), in the Bose-Marletto-Vedral (BMV) experiments, two particles are on the same horizontal line, so we can replace $\mathbf{x}$ with the scalar $x$. One can directly compute the density matrix element as:
@@
where $x_{1},x_{2},y_{1},y_{2}\in\{0,1\}$ representing the position of the corresponding particle. Since $\mathbf{x}_{1}=\pm\frac{a}{2}$ ($|{0}\rangle_{1}:=|{x_{1}=-\frac{a}{2}}\rangle$ and $|{1}\rangle_{1}:=|{x_{1}=+\frac{a}{2}}\rangle$), and $\mathbf{x}_{2}=\pm\frac{a}{2}+d$ ($|{0}\rangle_{2}:=|{x_{2}=d-\frac{a}{2}}\rangle$ and $|{1}\rangle_{2}:=|{x_{2}=d+\frac{a}{2}}\rangle$), after performing the partial transpose on particle 2 and diagonalizing the first-order correction in the zero eigenspace, the negativity of $\rho(\Delta t)$ is found to be
$\begin{split}\mathcal{N}(\rho^{T_{2}}(\Delta t))&=\frac{Gm^{2}\Delta t}{4\hbar}\Biggl[\bigl(\tilde{f}(0)-\tilde{f}(a)\bigr)+\\ &\frac{1}{2}\Bigl(\tilde{f}(d-a)+\tilde{f}(d+a)-2\tilde{f}(d)\Bigr)\Biggr],\end{split}$ (14)
where $\tilde{f}(z)=\text{erf}(z/(2\sigma))/z$ and $\tilde{f}(0)=1/(\sigma\sqrt{\pi})$. This total expression is manifestly non‑negative for all parameters. If we shift the coordinate of particle 2 by $a$, i.e., let $d\to d+a$, then the expression reduces to the one reported in Ref.[26], with distances $d$, $d+a$, and $d+2a$. In the limit $d\to\infty$ the second bracket vanishes, leaving a constant positive term proportional to $\tilde{f}(0)-\tilde{f}(a)>0$, which reflects the long‑range spatial correlation of the hidden detectors in the DP model. Hence, entanglement is already generated at first order in $\Delta t$ by the monitoring process alone.
3. About the quantum hidden degree of freedom in the monitoring process. In the collapse-based model of gravity [26, 25, 1], the SEM (see Eq. II) is used to analyze the DP and CML models. This SME stems from the continuous measurement theory in which there is a hidden degree of freedom that can monitor the observable of interest [25]. In the DP case, this hidden degree of freedom (also called ‘hidden detector’) weakly measures the mass density. It is well known that there is no consistent quantum-classical dynamics for quantum measurement problems [7, 9]. Solving the measurement problem requires either introducing non-commutative degrees of freedom into classical degrees of freedom (which contradicts the definition of classical systems) or the introduction of superselection rules [24, 12]. Indeed, continuous measurement theory is also introduced to address the quantum measurement problem, and it introduces a quantum hidden detector [27]. Thus, this hidden degree of freedom must be quantum-like. If the hidden degree of freedom is classical, it can not be used to monitor the observables of quantum systems. In [25], the author made the following statement:
@@
If the monitoring process of hidden detectors is not a physical process but merely a mathematical construct [25, 26, 1], collapse-based models would violate causality and the principle of locality, as they implicitly assume instantaneous interactions between masses. This assumption further leads to the paradoxical conclusion that all matter would spontaneously decohere and (weakly) entangle together.